OpenStudy (anonymous):

19. Solve the following by elimination: 3y - 1.6x = -15.4 8y - 2x = -32

OpenStudy (anonymous):

show your work please & thank you <3 I will give a medal (:

OpenStudy (anonymous):

you will hv to multiplu first bu 2 n 2nd by 1.6

OpenStudy (anonymous):

so it would be 3y - 1.6x * 2 = -15.4?

OpenStudy (anonymous):

multiply by*

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2(3y-1.6x) = 2*15.4

OpenStudy (anonymous):

oh okay, hold on let me solve that

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is it 2 * -15.4? at the end of the problem you gave me

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so would it be 6y - 3.2x = -30.8 ?

OpenStudy (anonymous):

yes now multiply 2nd by 1.6

OpenStudy (anonymous):

2nd?

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

8y - 2x = -32 this one

OpenStudy (anonymous):

so would it be 12,8y - 3.2x = -51.2

OpenStudy (anonymous):

good

OpenStudy (anonymous):

now,you have two equations 6y - 3.2x = -30.8 and 12.8y - 3.2x = -51.2 now if you sutract equation 2 from 1,-3.2x term will get canceled

OpenStudy (anonymous):

so you have (6y-3.2x) - (12.8y-3.2x) = -30.8 -(-51.2)

OpenStudy (anonymous):

so 6y-12.8y = =20.4

OpenStudy (anonymous):

-6.8y= 20.4 so y = -3

OpenStudy (anonymous):

put this value in other equation and find x

OpenStudy (anonymous):

3,2x - 3.2x ?

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

= 0 ?

OpenStudy (anonymous):

what do you wanna know exactly

OpenStudy (anonymous):

8y - 2x = -32 8(-3) -2x = -32 -24-2x=-32 2x=8 x=4